OpenStudy (axel.caballero):

Roy exclaims that his quadratic with a discriminant of -9 has no real solutions. Roy then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Roy, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (amistre64):

im with Roy ... why bother :)

OpenStudy (axel.caballero):

Youre funny @amistre64 but seriously your a moderator now, k'mon. Can you help ?

OpenStudy (mathstudent55):

The discriminant is \(b^2 - 4ac\). Come up with an equation of the form \(ax^2 + bx + c = 0\) for which \(b^2 - 4ac\) comes out negative.

OpenStudy (amistre64):

if Roy knows about a discriminant, then they most likely know the quadratic formula as well. Right?

OpenStudy (amistre64):

..and i know math55 here is capable

OpenStudy (mathstudent55):

Use a small number for b, and positive numbers for a and c, so that when you multiply 4ac you get a larger number than b^2.

OpenStudy (mathstudent55):

@axel.caballero Are you following along?

OpenStudy (axel.caballero):

@mathstudent55 yes, its just i only get on here and hour a day. what are the steps in solving this ?

OpenStudy (axel.caballero):

@amistre64 you should totes help me, this guy went offline.

OpenStudy (mathstudent55):

Let's come up with an example of a quadratic equation and check its discriminant. Let's try \( x^2 + 2x + 4\). In this case, \( a = 1, ~b = 2, ~c = 4\). The discriminant is \(b^2 - 4ac = 2^2 - 4 \cdot 1 \cdot 4= 4 - 16 = -12\) As you can see, the discriminant is negative, so there are no real solutions to this quadratic equation.

OpenStudy (axel.caballero):

Oh i see, this makes more sence. thanks! @mathstudent55